OpenStudy (anonymous):

help me plzzz Solve for y. a=(x+y+z)/3

OpenStudy (anonymous):

is this multiple choice?

OpenStudy (anonymous):

y = a - 3x - z y = a - x - 3z y = 3a - x - z

OpenStudy (anonymous):

y=3a-x-z